OpenStudy (anonymous):

Need help with chain rule please, especially when it comes to using it in conjunction with quotient rule.

OpenStudy (anonymous):

sorry typed that wrong y does \[\frac{ 3x-5 }{ (x-4)^{7} }\] become \[\frac{ 35-21x }{ (x-4)^{8} }+\frac{ 3 }{ (x-4)^{7} }\]

OpenStudy (anonymous):

the general form of \[F \prime (G(x))+G(x)\] I understand

OpenStudy (anonymous):

F′(G(x))+G'(x)

OpenStudy (anonymous):

If anyone could provide an explanation or pointers as to how I go about using the chain rule for cases such as the first example where I have \[\frac{ f(x) }{ g(x) }\]